# Verifiable Natural Language to Linear Temporal Logic Translation: A Benchmark Dataset and Evaluation Suite

### v1

@@ Abstract
Empirical evaluation of state-of-the-art natural-language (NL) to temporal-logic (TL) translation systems reveals near-perfect performance on existing benchmarks. However, current studies measure only the accuracy of the translation of NL logic into formal TL, ignoring a system’s capacity to ground atomic propositions into new scenarios or environments. This is a critical feature, necessary for the verification of resulting formulas in a concrete state space. Consequently, most NL-to-TL translation frameworks propose their own bespoke dataset in which the correct grounding is known a-priori, inflating performance metrics and neglecting the need for extensible, domain-general systems. In this paper, we introduce Verifiable Linear Temporal Logic Benchmark (VLTL-Bench), a unifying benchmark that measures verification and verifiability of automated NL-to-LTL translation. The dataset consists of three unique state spaces and thousands of diverse natural language specifications and corresponding formal specifications in temporal logic. Moreover, the benchmark contains sample traces to validate the temporal logic expressions. While the benchmark directly supports end-to-end evaluation, we observe that many frameworks decompose the process into i) lifting, ii) grounding, iii) translation, and iv) verification. The benchmark provides ground truths after each of these steps to enable researches to improve and evaluate different substeps of the overall problem. Using the benchmark, we evaluate several state-of-the-art NL-to-TL translation models and frameworks, including nl2spec, NL2TL, NL2LTL, sequence-to-sequence translation, and various prompting techniques. Our evaluation reveals that existing work is capable of reliably performing lifting and translation with high accuracy but struggle with grounding the translation in a state space, which stems from the lack of datasets to drive the research. Moreover, we provide traces that can be used for integrating verification into translation process, unlocking further directions for future research. To encourage methodologically sound advances in verifiable NL-to-LTL translation approaches, we release VLTL-Bench here: https://www.kaggle.com/datasets/dubascudes/vltl-bench.
@@ 1 Introduction
Formal verification is essential for the safe deployment of autonomous robots [24, 23], cyber–physical controllers [14], and safety-critical software systems [2]. Verification first begins with a specification that defines intent in precise temporal-logic (TL) [29, 3]. However, human stakeholders typically articulate intent in ambiguous natural language (NL) [26, 16, 15] and the conversion of this NL to TL is a challenging and time-consuming process that requires human experts [31, 4, 25]. Due to this complexity, automated translation of NL to TL has emerged as a core research problem [6, 33, 12, 28]. Recently, neural seq-to-seq models [11, 20, 13], grammar-constrained decoders [21, 9], and large language models (LLMs) [30, 6, 8, 7] have all demonstrated promising results on benchmark corpora, with reported string-exact accuracies often exceeding $90\%$. Despite these gains, we find existing evaluations provide an incomplete — and sometimes misleading — picture of true semantic competence. Specifically, we observe that high translation accuracies typically report the correctness of lifted (Section 2.1) TL predictions. While these expressions are syntactically correct, and indeed are equivalent to the expected lifted TL, there is little investigation into a framework’s ability to ground the lifted atomic propositions into a world state space. This oversight can be attributed to the widespread lack of an accompanying concrete world state space in existing corpora.
Figure 1: Overview of our dataset synthesis and evaluation pipline for NL to LTL translation. The scenario configuration is used to construct grounded APs as concrete action/target combinations in a given scenario. We generate unique NL representations from diverse references to the concrete variables. These free-form and concrete AP representations are inserted into our sentence, LTL, and trace templates to generate each corpora. Each component of our datasets can be used individually or combined to facilitate the holistic evaluation on NL-to-LTL translation frameworks.
Table 1: Comparison of existing LTL benchmarks and VLTL-Bench. Our datasets exhibit full coverage of logical/linguistic complexity, scenario inclusion, and verification support. $\times$ indicates total lack, $\sim$ indicates a feature could be estimated with processing, and $\checkmark$ indicates representation.
Dataset Complexity Inclusion
Linguistic Logical Lifting System Model Grounding Verification
CW (MacGlashan et al. [18]) $\times$ $\sim$ $\sim$ $\times$ $\times$
GLTL (Gopalan et al. [10]) $\times$ $\sim$ $\sim$ $\times$ $\times$
Navi (Wang et al. [27]) $\sim$ $\times$ $\times$ $\times$
Conformal (Wang et al. [28]) $\sim$ $\sim$ $\times$
VLTL-Bench Warehouse
VLTL-Bench Traffic Light
VLTL-Bench Search and Rescue
Current approaches for automated NL-to-TL translation are characterized by the use of multiple components, each with a distinct purpose in the end-to-end translation process. As noted earlier, we categorize these components into either lifting, grounding, or translation. Many frameworks employ different combinations of these components, such as nl2spec [7], which presupposes a pre-lifted and pre-grounded natural language specification as input. For example, an input "Globally, grant 0 and grant 1 do not hold at the same time until it is allowed," is assumed to be accompanied by a dictionary that maps the natural language sub-string "grant 1" to a known atomic proposition $g_{1}$. Therefore, this framework lacks the components for lifting and grounding, but includes a human auditing feature, which we omit from our analysis of automated translation frameworks. Meanwhile, another SOTA approach NL2TL [6] is divided into two processes. First, an LLM - in this case GPT3.5 [1] - is prompted to extract the atomic propositions (APs) from the NL input, creating “lifted NL" (NL without complex propositions). This new, less complex representation is then translated by a trained sequence-to-sequence model, such as T5 [22]. For example, an NL specification for a system: “You have an apple until you eat it.", will be converted into the lifted NL: “prop1 until prop2.". Here, prop1 is “You have an apple" and prop2 is “you eat it". This lifted NL is then translated to the lifted TL: “prop1 $\cup$ prop2." and the lifted propositions would be replaced.
These two frameworks are representative of the general research effort to produce better NL-to-TL translation frameworks, and yet no current frameworks have a holistic and faithful approach to translation. Therefore, we note the need for a benchmark that includes the data required to evaluate all components of the end-to-end translation pipeline - lifting, grounding, translation, and verification - in order to provide a complete picture of a framework’s true performance. We compare existing datasets and benchmarks on the basis of the linguistic and logical complexity of the end-to-end input and output, as well as the inclusion of features required to measure the performance of each individual component in Table 1.
We address these gaps with the Verifiable Linear Temporal Logic Benchmark (VLTL-Bench), a benchmark that grounds linear temporal logic (LTL) in a concrete world state space while broadening linguistic and logical coverage. VLTL-Bench introduces three new datasets which contain both raw and lifted versions of NL specifications, an AP-to-Condition dictionary, lifted and grounded LTL, and examples of satisfying and unsatisfying traces. The provision of each of these elements enables us to evaluate the four abstract components present in emerging NL-to-TL translation frameworks. Furthermore, the inclusion of example traces allows us to further evaluate the semantic consistency of translation and grounding attempts, resulting in a new verification accuracy performance metric, which measures whether an output translation holds on an a given trace. This feature permits consistency checking against real formal traces- a heretofore omitted feature of NL-to-TL translation corpora which will improve future explorations of NL-to-TL translation. In plausible real-world applications, a system may be given some example traces with which to grade its own translation attempts. This provides the opportunity to improve translation attempts without human auditing. We believe that the ability to measure the performance of individual components of a given framework will clarify the weaknesses of existing frameworks and guide future work. We outline our dataset synthesis and translation evaluation pipelines in Figure 1.
In summary, our contributions are: (i) a single, extensible benchmark and data synthesis pipeline for the evaluation of all current NL-to-TL translation components; (ii) a first-of-its-kind evaluation of the verification potential of existing frameworks, using our example satisfying and unsatisfying traces; and (iii) an empirical study that exposes both new failure modes in current translation methods and the severe accuracy decline when grounding is required. By providing a rigorous and extensible yardstick, VLTL-Bench can catalyze methodologically sound advances in NL-to-TL translation and, more broadly, in grounding large language models for formal reasoning tasks.
@@ 2.1 Preliminaries
In this section, we formally define a number of key terms necessary to describe and evaluate natural language-temporal logic translation systems. In particular, we have found that several terms are often conflated or co-defined in prior work, often leading to confusion regarding framework abilities. In order to provide a cogent description of these systems, as well as a robust evaluation, we define these terms as follows:
  1. Scenario: Referred to in existing work as the “World", “Environment", or “Space". A set $S$ of conditions appearing on a trace.
@@
2.2 Linear Temporal Logic
The syntax of LTL is given by the following grammar:
2.3 Benchmark Datasets
The Cleanup World (CW) dataset MacGlashan et al. [18] pairs NL commands with reward functions in an OO-MDP. While this work introduces a supervised grounding approach, the dataset used for training and evaluation is limited in terms of logical and linguistic complexity. This is due to the minimal set of APs used to construct the examples, and lack of diversity in both AP representation and sentence structure. Similarly, GLTL Gopalan et al. [10] gathers English sentences mapped to Geometric LTL formulas in robot-navigation scenarios, emphasizing stochastic-window operators but remaining modest in scale. The contents are remarkably similar to Cleanup World in both quality and metrics, with both datasets having the same number of unique temporal logic phrases and unique APs. The use of LLMs to enhance sentence diversity had not yet been adopted in the field at this time, and no alternate AP representations are introduced in the natural language. Therefore, the evaluation abilities of the GLTL dataset are limited in largely the same manner as Cleanup World. The Navi. corpus, introduced by Wang et al. [27] couples NL commands with LTL formulas and execution traces in a grid world. Its 226 unique APs make it a strong test of translation and lexical robustness, though this improvement comes at the cost of grounding- the Navi corpus does not specify formal APs, providing instead POS-tagged natural language representations, as shown below. Most recently, the ConformalNL2LTL dataset, introduced by Wang et al. [28], frames translation as open-vocabulary QA wrapped in conformal prediction (CP) to provide statistical guarantees. If the cumulative CP $p$-value falls below a user-specified threshold, the system escalates to a human; otherwise, it commits the subformula. The authors release a 1000-example corpus of grounded robot tasks partitioned by difficulty. We provide examples and a quantitative evaluation of existing datasets in Appendix A.3.
3 Verifiable Linear Temporal Logic Benchmark
Our framework is built around two core stages: Grounded Scenario Configuration, which defines the domain in which every proposition lives, and Data Synthesis, which instantiates our library of expert templates with scenario-specific atoms and then prunes any invalid NL. In the first stage, we provide the canonical formalization of the agent and the world it operates in. In the second stage, we apply the canonical APs obtained from the first to our expert-defined NL, TL, and trace templates.
@@ 3.1 Grounded Scenario Configuration
Each dataset is parameterized by a scenario—a small, declarative world model that provides:
  • Targets $\mathcal{L}$: e.g. {location, item} or {street, light}.
  • Actions $\mathcal{A}_{\mathrm{args}}$: verbs the agent may perform (e.g. idle, deliver(item, location), search(item)).
  • Target types $t\in\mathcal{T}$: each with its own action set $\mathcal{A}_{t}$ and argument signatures (e.g. deliver(item,location), avoid(threat)).
By swapping the scenario description of $\mathcal{L}$, $\mathcal{A}_{\mathrm{args}}$, and $\{\mathcal{A}_{t}\}$, the same logical and linguistic templates can be deployed in multiple domains without changing the core synthesis code.
@@ 3.2 Data Synthesis
To produce our datasets, we started with the 36 expert-crafted lifted NL-LTL pairs introduced by Cosler et al. [7] in the nl2spec benchmark and we crafted 7 of our own. We then crafted 43 pairs of corresponding example traces to each NL-LTL pair. Following this, we transform these lifted examples into a set of templates that can be used for diverse NL-LTL synthesis, introducing three substantial increases in complexity.
@@
    We generate more diverse words for objects, actions, and items as a potential scenario to enhance each example.
@@
    Each LTL skeleton (e.g. globally(prop_1 implies prop_2)) is coupled to action and target references from these scenarios.
@@
    Each example is given a corresponding pair of trace templates that yield a minimal satisfying ($\sigma_{good}\models\varphi_{G}$) and violating ($\,\sigma_{bad}\not\models\varphi_{G}$) trace.
To generate one example:
  1. 1.
    Atom sampling. Draw $k$ unique atomic propositions by randomly selecting actions and arguments from the scenario’s $\mathcal{A}_{\mathrm{ego}}\cup\mathcal{A}_{t}$ and $\mathcal{L}$.
  2. 2.
    Template instantiation. Uniformly choose a lifted template whose arity $\leq k$. Fill its LTL skeleton with the $k$ canonical atoms to obtain the grounded formula $\varphi_{g}$, and replace each atom by $\mathsf{prop}_{i}$ to obtain the lifted formula $\varphi_{L}$.
  3. 3.
    NL realization. Fill the semantic pattern with each atom’s surface-form (including articles/prepositions), apply morphological fixes (gerunds, capitalization), and record token-level spans. Emit both the free-form sentence and its grounded_sentence with explicit $\mathsf{prop}_{i}$ placeholders.
@@
    Trace filling. Apply the template’s trace builder to the list $[\mathsf{prop}_{1},\dots,\mathsf{prop}_{k}]$, yielding one good trace (satisfies $\varphi_{G}$) and one bad trace (violates $\varphi_{G}$).
  5. 5.
    Pruning. Submit each candidate NL sentence to GPT-4o for a lightweight check of the natural language components, removing any entries that the model flags as nonsensical or incorrect from the benchmark set.

### v2

@@ Abstract
Empirical evaluation of state-of-the-art natural language (NL) to temporal logic (TL) translation systems reveals near-perfect performance on existing benchmarks. However, current studies only measure the accuracy of the translation of NL logic into formal TL, ignoring a system’s capacity to ground atomic propositions into new scenarios or environments. This is a critical feature, necessary for the verification of resulting formulas in a concrete state space. In this paper, we introduce the Verifiable Linear Temporal Logic Benchmark (VLTL-Bench), a unifying benchmark for automated NL-to-LTL translation. The dataset consists of three unique state spaces and thousands of diverse natural language specifications and their corresponding formal temporal logic specifications. Moreover, the benchmark contains sample traces to verify the temporal logic expressions. While the benchmark directly supports end-to-end evaluation, we observe that many frameworks decompose the process into i) lifting, ii) grounding, iii) translation, and iv) verification. The benchmark provides ground truths after each of these steps to enable researchers to improve and evaluate different substeps of the overall problem. Using the benchmark, we evaluate several state‑of‑the‑art NL-to-TL translation models and frameworks, including nl2spec, NL2TL, NL2LTL, Lang2LTL, sequence-to-sequence translation, and various LLM prompting techniques. Our evaluation confirms that existing work is capable of reliably performing lifting and translation with high accuracy, while it exposes their struggles to ground the translation into a state space, which stems from the lack of existing datasets.
@@ 1 Introduction
Formal verification is essential for the safe deployment of autonomous robots (RobotLanguage; Spec2), cyber-physical controllers (SurveyTemporal), and safety-critical software systems (SAS_textbook_PoCPS). Verification first begins with a specification that defines intent in precise temporal logic (TL) (watson2005autonomous; bellini2000temporal). However, human stakeholders typically articulate intent in ambiguous natural language (NL) (veizaga2021systematically; lamar2009linguistic; lafi2021eliciting), and the conversion of this NL to TL is a challenging and time-consuming process that requires human experts (yin2024formal; cardoso2021review; thistle1986control). Due to this complexity, automated NL-to-TL translation has emerged as a core research problem (NL2TL; Natural2CTL_data; circuits_data; conformalLTL_data). Recently, neural sequence-to-sequence models (hahn2022formal; pan2023data; hsiung2022generalizing), grammar-constrained decoders (post2018fastlexicallyconstraineddecoding; geng2024grammarconstraineddecodingstructurednlp), and large language models (LLMs) (xu2024learning; NL2TL; NL2LTL; nl2spec) have all demonstrated promising results on benchmark corpora, with reported accuracies often exceeding $90\%$.
Despite these gains, evaluations are misleading as most datasets only test lifted translation, where temporal logic formulas contain abstract placeholders for atomic propositions (APs). The harder task of grounded translation—instantiating APs with domain-specific actions and arguments—is usually left unmeasured. This imbalance stems from limitations of current datasets, which omit the annotations required to separately evaluate lifting, translation, and grounding. As a result, current frameworks optimize for partial tasks, leaving open the more difficult but necessary problem of grounding for producing fully executable specifications.
Figure 1: Overview of our dataset synthesis and evaluation framework for NL-to-LTL translation. The framework used a configuration file to define concrete and unique scenarios. The data synthesis generates the NL and TL pairs with associated traces for verification while providing ground truth results for intermediate components.
Benchmarks for NL-to-TL translation include CW (CW_data), GLTL (GLTL-data), Navi (Navi_data), and Conformal (conformalLTL_data). Their limitations are fourfold. (i) Although recent frameworks decompose the task into lifting, translation, and grounding, these benchmarks supply ground truth only for the end-to-end result (NL-TL pairs), preventing assessment of intermediate components. (ii) CW and GLTL omit grounding entirely, yielding translations without executable semantics. For example, the NL specification: “Go to the green room and then go to the blue room.” is mapped to the LTL expression: $\diamondsuit G\rightarrow\diamondsuit B$”, without providing a grounded definition of the predicates $G$ and $B$. (iii) Navi and Conformal nominally support grounding but rely on overly simplistic state spaces (e.g., Navi’s colored-room grid), which fails to capture the referential and contextual ambiguities of natural language. (iv) Execution traces/trajectories for independent semantic verification (e.g., via model checking), are not provided, preventing rigorous evaluation.
In this paper, we introduce the Verifiable Linear Temporal Logic Benchmark (VLTL-Bench), a benchmark that grounds linear temporal logic (LTL) in a concrete world state space while broadening linguistic and logical coverage through more diverse atomic propositions. As illustrated in Figure 1, VLTL-Bench exposes every stage of the NL-to-TL pipeline: raw and lifted NL specifications, an AP-to-Reference dictionary, lifted and grounded LTL formulas, and both satisfying and unsatisfying traces. Our dataset synthesis and evaluation framework for NL-to-LTL translation leverages scenario configurations to construct grounded action/target combinations, from which we synthesize diverse natural language representations and integrate them into sentence, LTL, and trace templates, yielding corpora whose components can be used individually or combined for holistic evaluation. This layered design makes it possible to isolate performance on lifting, translation, grounding, and verification individually, while also enabling end-to-end evaluation. We provide four scenario configuration files and construct a Kitchen Assistant, Traffic light, Search & Rescue, and Warehouse dataset. Using these four datasets we evaluate the capabilities and limitations of state-of-the-art NL to TL translation frameworks. In summary, we propose: (i) a single, extensible benchmark for evaluating all NL-to-TL translation components; (ii) the first verification evaluation using satisfying and unsatisfying traces; and (iii) an empirical study that reveals both new failure modes in current methods and the severe accuracy decline when grounding is required.
The remainder of this paper is organized as follows. Section 2 covers preliminaries for LTL and model checking. Section 3 contains a detailed description of the Verifiable Linear Temporal Logic Benchmark datasets, as well as details on how they were synthesized. Section 4 includes an evaluation of current NL-to-TL frameworks on both Verifiable Linear Temporal Logic Benchmark and existing datasets. We conclude our paper in 5. Additional details may be found in the Appendix A
@@ 2 Background and Related Work
In this section we introduce necessary notation and background information on temporal logic systems including terminology, linear temporal logic symbols, and existing NL-to-TL datasets.
Linear Temporal Logic. The syntax of LTL is given by the following grammar:
@@ 2.1 Preliminaries
In this section, we formally define a number of key terms necessary to describe and evaluate NL-to-TL translation systems. In order to provide a cogent description of these systems, as well as a robust evaluation, we define these terms as follows:
  1. Scenario: Referred to in existing work as the “World”, “Environment”, or “Space”. A set $S$ of conditions appearing on a trace.
@@
Table 1: Comparison of existing LTL benchmarks and VLTL-Bench. We report the number of unique words across all NL specifications and the number of unique LTL specifications. Additionally, we report support for lifting, grounding, and verification.
Dataset # Words # TL Lifting Translation Grounding Verification
CW (CW_data) 184 37 $\sim$ $\times$ $\times$
GLTL (GLTL-data) 183 37 $\sim$ $\times$ $\times$
Navi (Navi_data) 131 6414 $\times$ $\sim$ $\times$
Conformal (conformalLTL_data) 439 212 $\sim$ $\checkmark$ $\sim$ $\times$
VLTL-Bench Warehouse 1028 5991
VLTL-Bench Traffic Light 217 6196
VLTL-Bench Search and Rescue 245 5425
VLTL-Bench Kitchen Assistant 376 7385
2.2 Existing Benchmark Datasets
In this section, we review existing benchmarks for NL-to-TL translation. We compare these corpora in terms of linguistic and logical complexity, and support for evaluation of different framework modules in Table 2.1. We measure the complexity using the number of unique words appearing in natural language specifications (#Words), as well as the number of unique temporal logic expressions (#TL). In terms of modules, we report if a dataset has support for evaluation of lifting, grounding, and verification. We also provide examples from existing datasets in Appendix A.5.
In Table 2.1, we observe that the older datasets Cleanup World (CW) (CW_data) and GLTL (GLTL-data) from the pre-LLM era have limited complexity both in terms of unique words and temporal logic expressions. While they support evaluation of translation, the lifting data is not explicitly given, the APs do not vary in their form to any meaningful degree, and they can be lexically identified in both the NL and TL elements of each entry ("green room" $\leftrightarrow G$, "blue room" $\leftrightarrow B$, etc.). The Navi. corpus, introduced by (Navi_data), couples NL commands with LTL formulas in a grid world. As Table 2.1 shows, Navi exhibits a substantial increase in logical complexity, with 6,414 unique formulas and support for partial grounding. Its 221 unique APs make it a strong test of translation and lexical robustness, though this improvement comes at the cost of well-defined grounding rules: the corpus does not specify formal APs, providing instead POS-tagged natural language representations. As reflected in Table 2.1, the Conformal (conformalLTL_data) dataset introduces 439 unique words and 212 formulas with explicit grounding, but its scale is modest at 1,000 examples. In contrast, VLTL-Bench provides a testbed suited to holistic evaluation across lifting, translation, grounding, and verification. We provide a more detailed quantitative comparison between these datasets and VLTL-Bench in Section 3.4.
3 The Verifiable Linear Temporal Logic Benchmark
In the following subsections, we first introduce Grounded Scenario Configuration, which formalizes the world model by defining types, targets, and actions that ensure well-typed logical atoms. We then describe our Data Synthesis pipeline, which instantiates expert-crafted NL–LTL templates with scenario-specific atoms to produce paired sentences, formulas, and traces. Next, we present the Metrics used to evaluate each stage of the NL-to-LTL pipeline, and finally, we detail the Datasets generated from three scenario definitions, highlighting their unique challenges and properties.
@@ 3.1 Grounded Scenario Configuration
To formalize how natural language specifications map onto executable logical structures, we distinguish three interconnected components: types, targets, and actions. Types serve as abstract categories that describe what kinds of objects or entities an action can take as input (e.g., a location, an item, or a threat). Targets are the grounded instantiations of these action–type combinations, where abstract slots are filled with concrete constants. Actions are verbs that capture the capabilities of the agent; each action comes with a signature that specifies the expected types of its arguments. Together, this hierarchy ensures that linguistic expressions can be systematically mapped into well-typed logical atoms: types constrain argument structure, actions define the permissible predicates, and targets bind them to domain-specific instances. Each dataset is parameterized by a scenario—a small, declarative world model that provides:
Types $t\in\mathcal{T}$: denotes the sort of parameters accepted by an action (e.g. item or location).
Targets $\mathcal{L}$: Specific instances of typed arguments, (e.g. an argument apple of type item, or an argument loading_dock of type location).
Actions $\mathcal{A}_{\mathrm{args}}$: verbs the agent may perform, which may have one or more targets, (e.g. idle() has no targets, deliver(apple, loading_dock) takes two—item and location).
@@ 3.2 Data Synthesis
To produce our datasets, we began with the 36 expert-crafted lifted NL-LTL pairs of the nl2spec benchmark (nl2spec), and we added 7 new ones of our own (provided in Appendix A.4). We then transformed these 43 examples into templates to support diverse NL–LTL synthesis. Finally, for each NL–LTL example, we crafted one pair of traces—one satisfying and one violating.
Each dataset entry includes a tuple of these three artifacts,
[…]
and is algorithmically constructed with the following steps:
@@
    Template selection. Uniformly choose a lifted template. Each template has an arity that determines how many atomic propositions must be instantiated.
@@
    Atom sampling. For each argument slot in the template, draw a unique atomic proposition by randomly selecting actions and arguments from the scenario’s $\mathcal{A}_{t}$ and $\mathcal{L}$. Let $k$ denote the total number of sampled atoms. Fill the LTL skeleton with these $k$ atoms to obtain the grounded formula $\varphi_{G}$, and replace each atom by $\mathsf{prop}_{i}$ to obtain the lifted formula $\varphi_{L}$.
@@
    NL realization. Fill the template pattern with each atom’s surface form (including articles/prepositions), apply morphological fixes (gerunds, capitalization), and record token‐level spans. Emit both the free‐form sentence and its grounded_sentence with explicit $\mathsf{prop}_{i}$ placeholders.
@@
    Trace filling. Apply the template’s trace patterns to the list $\mathsf{prop}_{1},\dots,\mathsf{prop}_{k}]$, yielding one positive trace (satisfies $\varphi_{G}$) and one negative trace (violates $\varphi_{G}$).
This rich annotation supports four independent evaluation axes, displayed in Figure 2.